The raise in temperature of a given mass of an ideal gas at constant pressure and at temperature 27°C to double its volume is. (1) 327°C. (2) 54°C. (3) 300°C. (4) 600°C

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Anjali
 
Use 
PV /T= constant
Hence V1/T1 = V2/T2
V1/300 = 2 V1/T2
T2 = 600 k
T2 = 327
 
Regards
Arun (askIITians forum expert)